# Complete $k$-partite entanglement measure

### v3

@@ Abstract
Keywords: $k$-partite entanglement, $k$-partite entanglement measure, Unified/Complete $k$-partite entanglement measure
@@ 1 Introduction
In 2005, Gühne et al. introduced the $k$-partite entanglement in Ref. [1]. It is closely related to the $k$-producible state: if a quantum state is not $k$-producible, it is termed $(k+1)$-partite entangled. While the $k$-entanglement reflects how many splitted subsystems are entangled under partitions of the systems [2, 3], the $k$-partite entanglement concentrate on at most how many particles in the global system are entangled but separable from other particles [1, 4, 5]. It has been shown that $k$-producibility plays a crucial role in both quantum nonlocality [6, 7, 8, 9] and quantum metrology [10]. Particularly, $k$-producibly entangled states for larger $k$ exhibit higher sensitivity in phase estimation [11, 12, 13].
Clearly, these two ways of exhibiting entanglement, i.e., the $k$-entanglement and the $k$-partite entanglement, are complementary to each other in characterizing the multipartite entanglement which remains challenging to understand undeniably since the complexity increases substantially with the number of parties [14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Recently, the $k$-partite entanglement measure based on concurrence have been presnted [4, 5]. Very recently, we established the theory of the complete $k$-entanglement measure in Ref. [2]. It was shown that, in the framework of the complete measure of quantum correlation, the distribution of the correlation could be depicted exhaustively since the correlation could be compared not only between the global system and the subsystem (or the systems under arbitrary partition) but also between different subsystems (or the systems under arbitrary partition) [2, 21, 23, 24, 25, 26]. Along this line, the aim of this paper is to discuss how can we establish the axiomatic postulates for the complete $k$-partite entanglement measure.
The rest of the paper is arranged as follows. We review the concept of $k$-partite entanglement, the $k$-partite entanglement measures proposed in Ref. [4, 5], and the coarsening relation of multipartite partitions in Sec. 2. In Sec. 3, we present the definition of the complete $k$-partite entanglement measure, and then give two general ways of constructing $k$-partite entanglement measures and discuss whether they are complete in Sec. 4. Sec. 5 lists some examples of $k$-partite entanglement measures according to the methods in Sec. 4. Finally, in Sec. 6, we summarize the results of the paper.
@@ 2 Notations and Preliminaries
For convenience of discussing the complete measure of the $k$-partite entanglement in the next sections, we review some basic notations and terminologies in Sec 2.1, and introduce the $k$-partite entanglement measures proposed in literature so far in Sec. 2.2. We then introduce the coarsening relation of the multipartite partitions which is necessary when we discuss the completeness of a multipartite quantum correlation measure (also see in Ref. [23, 24]).
@@ 2.1 $k$-partite entanglement
A pure state $|\psi\rangle$ of an $n$-partite system $A_{1}A_{2}\cdots A_{n}$ with state space $\mathcal{H}^{A_{1}A_{2}\cdots A_{n}}$ is called $k$-producible ($1\leqslant k\leqslant n-1$), if it can be represented as [1]
@@
under some $k$-fineness partition $X_{1}|X_{2}|\cdots|X_{m}$ of $A_{1}A_{2}\cdots A_{n}$. Let $\mathcal{S}^{X}$ be the set of all density operators acting on the state space $\mathcal{H}^{X}$. For mixed state $\rho\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$, if it can be written as a convex combination of $k$-producible pure states, i.e., $\rho=\sum\limits_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$ with $|\psi_{i}\rangles$ are $k$-producible, it is called $k$-producible, where the pure state $|\psi_{i}\rangles$ might be $k$-producible in different $k$-fineness partitions. If a quantum state is not $k$-producible, it is termed $(k+1)$-partite entangled. Note that if $|\psi\rangle$ admits the form as in Eq. (1), it is called $m$-separable. $|\psi\rangle$ is $m$-entangled if it is not $m$-separable. An $n$-partite mixed state $\rho$ is $m$-separable if it can be written as $\rho=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$ with $|\psi_{i}\rangles$ are $m$-separable, wherein the contained $\{|\psi_{i}\rangle\}$ can be $m$-separable with respect to different $m$-partitions. Otherwise, it is called $m$-entangled. By definition, the $k$-partite entanglement is different from the $k$-entanglement in general, but they are equivalent only in some special cases. For example, the $n$-partite entangled state is just the genuine multipartite entangled state and the one-producible state coincides with the fully separable state. If $|\psi\rangle^{ABC}$ is a genuine entangled state, then $|\psi\rangle^{ABC}|\psi\rangle^{D}|\psi\rangle^{E}|\psi\rangle^{F}$ is four-separable and three-partite entangled state. Also note that, a state of which some reduced state of $m$ parties is genuinely entangled, contains $m$-partite entanglement, but not vice versa in general [1]. For more clarity, we compare $3$-partite entangled pure state with $3$-entangled pure state in Fig. 1.
@@
A pure state $|\phi\rangle$ is said to be genuinely $k$-producible (or genuinely $k$-partite entangled) [1] if it is $k$-producible but not (k1)(k-\!1)-producible. A mixed state $\rho\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$ is genuinely $k$-producible if it is $k$-producible and for any $k$-producible pure states ensemble of $\rho$, $\rho=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$, there is at least one $|\psi_{i}\rangle$ is genuinely $k$-producible.
@@ 2.2 $k$-partite entanglement measures via $q$-concurrence and $\alpha$-concurrence
Hong et al. presented a $k$-PEMo in Ref. [4] via concurrence. For any pure state $|\psi\rangle\in\mathcal{H}^{A_{1}A_{2}\cdots A_{n}}$, the $k$-PEMo was defined as [4]
@@
Very recently, Li et al. proposed two $k$-PEMos in Ref. [5]. For any pure state $|\psi\rangle\in\mathcal{H}^{A_{1}A_{2}\cdots A_{n}}$, the $k$-PEMo via the $q$-concurrence was defined as [5]
@@
and the $k$-PEMo via the $\alpha$-concurrence was expressed by [5]
@@
where the minimum is taken over all the (k1)(k\!-\!1)-fineness partitions in $\Gamma_{\!k-\!1}^{f}$. Note here that, the notations here are different from $E_{q-k}$ and $E_{\alpha-k}$ in Ref. [5] ($C_{q(k+1)}=E_{q-k}$, $C_{\alpha(k+1)}=E_{\alpha-k}$). Ref. [5] also gave the following two $k$-PEMos:
@@
where $\rho_{X_{t(i)}}$ is the reduced density operator with respect to subsystem $X_{t(i)}$}, and $m_{i}$ refers to $\gamma_{i}$ is a $m_{i}$-partition, $\left|\Gamma_{\!k-\!1}^{f}\right|$ is the cardinal number of $\Gamma_{\!k-\!1}^{f}$. The notations in Eqs. (6), (7) are different from $\varepsilon_{q-k}$ and $\varepsilon_{\alpha-k}$ in Ref. [5] ($C_{G,q(k+1)}=\varepsilon_{q-k}$, $C_{G,\alpha(k+1)}=\varepsilon_{\alpha-k}$).
@@ 2.3 Coarsening relation of multipartite partitions
Let $X_{1}|X_{2}|\cdots|X_{k}$ and $Y_{1}|Y_{2}|\cdots|Y_{l}$ be two partitions of $A_{1}A_{2}\cdots A_{n}$ or subsystem of $A_{1}A_{2}\cdots A_{n}$, $k\leq n$, $l\leq n$. We denote by [2, 23]
@@
For any subsystem of $A_{1}A_{2}\cdots A_{n}$ with arbitrary partition, it can always be derived from the global system via the coarsening relations (a)-(c) or some of them. So, based on these three coarsening relations, we can analyze not only the information relation between any subsystem and the global subsystem but also the information relation between any subsystems in detail. For instance, based on these coarsening relations, we have established the complete global entanglement measure [21, 24], the complete genuine entanglement measure [23, 24], the complete multipartite quantum discord [25], the complete multipartite quantum mutual information [26] and the complete $k$-entanglement measure [2]. Furthermore, we discussed the complete monogamy relation of these measures [2, 21, 23, 24, 25, 26], where exploring the monogamy relation of the quantum correlations is one of the fundamental tasks in the quantum resource theory [21, 25, 27, 28, 29, 30, 31, 32, 33, 34].
@@ 3 Completeness of the $k$-PEM
When we deal with the various quantum correlations living in a multipartite system, the most quintessential relation in a multipartite system is indeed the the coarsening relation, i.e., the coarsening relations of type (a)-(c). The “completeness” of a measure for multipartite quantum correlation mainly refers to that there is a unified criterion for quantifying different subsystems or systems under different partition, which means the amount of the quantum correlations contained in different particles or particles under arbitrary partition can be compared with each other consistently and compatibly [2, 21, 23, 25, 26]. It makes up for the previous bipartite measure which can only quantify the quantum correlation under the given bipartite splitting. By reviewing the key point in defining a complete measure of quantum correlation [2, 21, 23, 25, 26], we can conclude that there are two steps to reveal such a completeness of a given measure: the first step is the unification condition which is mainly related to the coarsening relation of type (a), and the second one is the hierarchy condition which is determined by the coarsening relation of type (b).
@@
[…]
@@
which is larger than $E_{(3)}(|\psi\rangle)$ whenever $|\psi\rangle^{DE}$, $|\psi\rangle^{GH}$ and $|\psi\rangle^{IJ}$ are entangled states. Namely, Eq. (13) is violated under the general coarsening operation of type (b). We thus adjust the hierarchy condition (i.e., the tight coarsening monotone condition) as Eq. (13) in stead of the “strict” hierarchy condition in Refs. [2, 21].
@@
for any $\rho\in\mathcal{S}^{ABCD}$.
@@ 4 Two classes of $k$-PEMos
[…]
@@ 4.1 $k$-PEMo from unified multipartite entanglement measure
[…]
@@
Proof.
Items (i) and (ii) are straightforward. For any $|\psi\rangle\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$, we suppose that $E_{(k)}(|\psi\rangle)=\sum\limits_{j=1}^{r}E^{(s(t_{j}))}(|\psi\rangle^{X_{t_{j}}})>0$ for some partition $X_{t_{1}}|X_{t_{2}}|\cdots|X_{t_{r}}|X_{*}$. It turns out that $E_{(k)}(|\psi\rangle^{X_{t_{j}}})=E_{(k-1)}(|\psi\rangle^{X_{t_{j}}})$ for any $j$, $E_{(k)}(|\phi\rangle^{X_{*}})=0$, but it is possible that $E_{(k-1)}(|\phi\rangle^{X_{*}})>0$, which implies (iii) is true. For any partition $X_{1}|X_{2}|\cdots|X_{p}$ of $A_{1}A_{2}\cdots A_{n}$, $p\geqslant k$, we consider $X_{2}|\cdots|X_{p}$ w.n.l.g., namely the partition that by discarding $X_{1}$ from $X_{1}|X_{2}|\cdots|X_{p}$. If $X_{1}$ is $X_{t_{j}}$ or some subsytem(ssome) of $X_{t_{j}}$, (iv) is clear since $E^{(s)}$ is unified [24] [a unified MEM is decreasing under the coarsening relation of type (a)]. If $X_{1}$ is $X_{*}$ or some subsytem(s) of $X_{*}$, (iv) is clear since $E_{(k)}(|\psi\rangle)=E_{(k)}(\rho^{X_{2}|\cdots|X_{p}})$. If $X_{1}$ cotains some subsytem(s) of some $X_{t_{j}}$, (iv) is also ture since $E^{(s)}$ is unified. The other cases can be argued similarly.
If the reduced function $h$ is subadditive, then $E^{(n)}$ is complete [21, 24], which means that $E^{(n)}$ is nonincreasing under the coarsening operation of type (b). This completes the proof. ∎
Another candidate for the unified global MEM is the one defined by the sum of all bipartite entanglement [22], i.e.,
@@
where $h$ is some given reduced function. We denote by $\mathcal{E}_{(k)}$ the quantity that is defined as in Eq. (20) just with $\mathcal{E}^{(s(t_{j}))}$ replacing $E^{(s(t_{j}))}$. Using similar arguments as in the proof of Theorem 1, we can conclude the following theorem.
Theorem 2.
@@
In Eq. (20), $|\psi\rangle^{X_{t_{j}}}$ is genuinely entangled, so $E^{(s(t_{j}))}$ can be any genuine entanglement measure instead. For example, the genuine entanglement measure from the minimal reduced function, which is defined by [24]
$\varepsilon_{g^{\prime\prime}}^{(n)}(|\psi\rangle^{A_{1}A_{2}\cdots A_{n}})=h(\rho^{X})$
where $h$ is some given reduced function, $X\subsetneq\{A_{1},A_{2},\dots,A_{n}\}$, namely the minimum runs over all possible reduced states. Then the corresponding $k$-PEMo is not unified in general since $\varepsilon_{g^{\prime\prime}}^{(n)}$ may increase under the coarsening relation of type (a) [24].

### v4

@@ 1 Introduction
In 2005, Gühne et al. introduced the $k$-partite entanglement in Ref. Guhne2005njp . It is closely related to the $k$-producible state: if a quantum state is not $k$-producible, it is termed $(k+1)$-partite entangled. While the $k$-entanglement reflects how many splitted subsystems are entangled under partitions of the systems Guo2024pra ; Hong2012pra , the $k$-partite entanglement concentrate on at most how many particles in the global system are entangled but separable from other particles Guhne2005njp ; Hong2023epjp ; Li2025aqt . It has been shown that $k$-producibility plays a crucial role in both quantum nonlocality Curchod2015pra ; Liang2015prl ; Lin2019pra ; Lu2018prx and quantum metrology Toth2014jpa . Particularly, $k$-producibly entangled states for larger $k$ exhibit higher sensitivity in phase estimation Hyllus2012pra ; Gessner2018prl ; Qin2019npj .
Clearly, these two ways of exhibiting entanglement, i.e., the $k$-entanglement and the $k$-partite entanglement, are complementary to each other in characterizing the multipartite entanglement which remains challenging to understand undeniably since the complexity increases substantially with the number of parties Horodecki2009 ; Verstraete2003pra ; Gour2010prl ; Jungnitsch2011prl ; Xie2021prl ; Jacob2021prl ; Li2024pra ; Guo2020pra ; Guo2022jpa ; Guo2022entropy . Recently, the $k$-partite entanglement measure based on concurrence have been presnted Hong2023epjp ; Li2025aqt . Very recently, we established the theory of the complete $k$-entanglement measure in Ref. Guo2024pra . It was shown that, in the framework of the complete measure of quantum correlation, the distribution of the correlation could be depicted exhaustively since the correlation could be compared not only between the global system and the subsystem (or the systems under arbitrary partition) but also between different subsystems (or the systems under arbitrary partition) Guo2024pra ; Guo2020pra ; Guo2022entropy ; Guo2024rip ; G2021qst ; Guo2023pra . Along this line, the aim of this paper is to discuss how can we establish the axiomatic postulates for the complete $k$-partite entanglement measure.
The rest of the paper is arranged as follows. We review the concept of $k$-partite entanglement, the $k$-partite entanglement measures proposed in Ref. Hong2023epjp ; Li2025aqt , and the coarsening relation of multipartite partitions in Sec. 2. In Sec. 3, we present the definition of the complete $k$-partite entanglement measure, and then give two general ways of constructing $k$-partite entanglement measures and discuss whether they are complete in Sec. 4. Sec. 5 lists some examples of $k$-partite entanglement measures according to the methods in Sec. 4. Finally, in Sec. 6, we summarize the results of the paper.
@@ 2 Notations and Preliminaries
For convenience of discussing the complete measure of the $k$-partite entanglement in the next sections, we review some basic notations and terminologies in Sec 2.1, and introduce the $k$-partite entanglement measures proposed in literature so far in Sec. 2.2. We then introduce the coarsening relation of the multipartite partitions which is necessary when we discuss the completeness of a multipartite quantum correlation measure (also see in Ref. Guo2022entropy ; Guo2024rip ).
@@ 2.1 $k$-partite entanglement
A pure state $|\psi\rangle$ of an $n$-partite system $A_{1}A_{2}\cdots A_{n}$ with state space $\mathcal{H}^{A_{1}A_{2}\cdots A_{n}}$ is called $k$-producible ($1\leqslant k\leqslant n-1$), if it can be represented as Guhne2005njp
@@
under some $k$-fineness partition $X_{1}|X_{2}|\cdots|X_{m}$ of $A_{1}A_{2}\cdots A_{n}$. Let $\mathcal{S}^{X}$ be the set of all density operators acting on the state space $\mathcal{H}^{X}$. For mixed state $\rho\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$, if it can be written as a convex combination of $k$-producible pure states, i.e., $\rho=\sum\limits_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$ with $|\psi_{i}\rangles$ are $k$-producible, it is called $k$-producible, where the pure state $|\psi_{i}\rangles$ might be $k$-producible in different $k$-fineness partitions. If a quantum state is not $k$-producible, it is termed $(k+1)$-partite entangled. Note that if $|\psi\rangle$ admits the form as in Eq. (1), it is called $m$-separable. $|\psi\rangle$ is $m$-entangled if it is not $m$-separable. An $n$-partite mixed state $\rho$ is $m$-separable if it can be written as $\rho=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$ with $|\psi_{i}\rangles$ are $m$-separable, wherein the contained $\{|\psi_{i}\rangle\}$ can be $m$-separable with respect to different $m$-partitions. Otherwise, it is called $m$-entangled. By definition, the $k$-partite entanglement is different from the $k$-entanglement in general, but they are equivalent only in some special cases. For example, the $n$-partite entangled state is just the genuine multipartite entangled state and the one-producible state coincides with the fully separable state. If $|\psi\rangle^{ABC}$ is a genuine entangled state, then $|\psi\rangle^{ABC}|\psi\rangle^{D}|\psi\rangle^{E}|\psi\rangle^{F}$ is four-separable and three-partite entangled state. Also note that, a state of which some reduced state of $m$ parties is genuinely entangled, contains $m$-partite entanglement, but not vice versa in general Guhne2005njp . For more clarity, we compare $3$-partite entangled pure state with $3$-entangled pure state in Fig. 1.
@@
A pure state $|\phi\rangle$ is said to be genuinely $k$-producible (or genuinely $k$-partite entangled) Guhne2005njp if it is $k$-producible but not (k1)(k-\!1)-producible. A mixed state $\rho\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$ is genuinely $k$-producible if it is $k$-producible and for any $k$-producible pure states ensemble of $\rho$, $\rho=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$, there is at least one $|\psi_{i}\rangle$ is genuinely $k$-producible.
@@ 2.2 $k$-partite entanglement measures via $q$-concurrence and $\alpha$-concurrence
Hong et al. presented a $k$-PEMo in Ref. Hong2023epjp via concurrence. For any pure state $|\psi\rangle\in\mathcal{H}^{A_{1}A_{2}\cdots A_{n}}$, the $k$-PEMo was defined as Hong2023epjp
@@
Very recently, Li et al. proposed two $k$-PEMos in Ref. Li2025aqt . For any pure state $|\psi\rangle\in\mathcal{H}^{A_{1}A_{2}\cdots A_{n}}$, the $k$-PEMo via the $q$-concurrence was defined as Li2025aqt
@@
and the $k$-PEMo via the $\alpha$-concurrence was expressed by Li2025aqt
@@
where the minimum is taken over all the (k1)(k\!-\!1)-fineness partitions in $\Gamma_{\!k-\!1}^{f}$. Note here that, the notations here are different from $E_{q-k}$ and $E_{\alpha-k}$ in Ref. Li2025aqt ($C_{q(k+1)}=E_{q-k}$, $C_{\alpha(k+1)}=E_{\alpha-k}$). Ref. Li2025aqt also gave the following two $k$-PEMos:
@@
where $\rho_{X_{t(i)}}$ is the reduced density operator with respect to subsystem $X_{t(i)}$}, and $m_{i}$ refers to $\gamma_{i}$ is a $m_{i}$-partition, $\left|\Gamma_{\!k-\!1}^{f}\right|$ is the cardinal number of $\Gamma_{\!k-\!1}^{f}$. The notations in Eqs. (6), (7) are different from $\varepsilon_{q-k}$ and $\varepsilon_{\alpha-k}$ in Ref. Li2025aqt ($C_{G,q(k+1)}=\varepsilon_{q-k}$, $C_{G,\alpha(k+1)}=\varepsilon_{\alpha-k}$).
@@ 2.3 Coarsening relation of multipartite partitions
Let $X_{1}|X_{2}|\cdots|X_{k}$ and $Y_{1}|Y_{2}|\cdots|Y_{l}$ be two partitions of $A_{1}A_{2}\cdots A_{n}$ or subsystem of $A_{1}A_{2}\cdots A_{n}$, $k\leq n$, $l\leq n$. We denote by Guo2024pra ; Guo2022entropy
@@
For any subsystem of $A_{1}A_{2}\cdots A_{n}$ with arbitrary partition, it can always be derived from the global system via the coarsening relations (a)-(c) or some of them. So, based on these three coarsening relations, we can analyze not only the information relation between any subsystem and the global subsystem but also the information relation between any subsystems in detail. For instance, based on these coarsening relations, we have established the complete global entanglement measure Guo2020pra ; Guo2024rip , the complete genuine entanglement measure Guo2022entropy ; Guo2024rip , the complete multipartite quantum discord G2021qst , the complete multipartite quantum mutual information Guo2023pra and the complete $k$-entanglement measure Guo2024pra . Furthermore, we discussed the complete monogamy relation of these measures Guo2024pra ; Guo2020pra ; Guo2022entropy ; Guo2024rip ; G2021qst ; Guo2023pra , where exploring the monogamy relation of the quantum correlations is one of the fundamental tasks in the quantum resource theory Guo2020pra ; G2021qst ; Coffman ; Pawlowski ; streltsov2012are ; Augusiak2014pra ; Osborne ; Dhar ; GG ; GG2019 .
@@ 3 Completeness of the $k$-PEM
When we deal with the various quantum correlations living in a multipartite system, the most quintessential relation in a multipartite system is indeed the the coarsening relation, i.e., the coarsening relations of type (a)-(c). The “completeness” of a measure for multipartite quantum correlation mainly refers to that there is a unified criterion for quantifying different subsystems or systems under different partition, which means the amount of the quantum correlations contained in different particles or particles under arbitrary partition can be compared with each other consistently and compatibly Guo2024pra ; Guo2020pra ; Guo2022entropy ; G2021qst ; Guo2023pra . It makes up for the previous bipartite measure which can only quantify the quantum correlation under the given bipartite splitting. By reviewing the key point in defining a complete measure of quantum correlation Guo2024pra ; Guo2020pra ; Guo2022entropy ; G2021qst ; Guo2023pra , we can conclude that there are two steps to reveal such a completeness of a given measure: the first step is the unification condition which is mainly related to the coarsening relation of type (a), and the second one is the hierarchy condition which is determined by the coarsening relation of type (b).
@@
holds for all states $\rho\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$ whenever $X_{1}|X_{2}|\cdots|X_{p}\succ^{a}Y_{1}|Y_{2}|\cdots|Y_{q}$ with $k\leqslant q\leqslant p$. Item (i) is clear, i.e., the symmetry is an inherent feature of entanglement measure indeed. The $k$-partite entanglement contained in $A_{1}A_{2}\cdots A_{r}\otimes A_{r+1}A_{r+2}\cdots A_{n}$ is composed of two parts, i.e., $A_{1}A_{2}\cdots A_{r}$ and $A_{r+1}A_{r+2}\cdots A_{n}$. So we demand item (ii). If a state is (k1)(k-1)-producible, it must be $k$-producible, but not vice versa, so we require condition (iii). For the generalized $n$-qudit GHZ state $\frac{1}{\sqrt{d}}(|00\cdots 0\rangle+|11\cdots 1\rangle+\cdots|d-1\rangle|d-1\rangle\cdots|d-1\rangle)$, Eq. (12) is always true for any $k$-PEM. In addition, $E_{(k)}(|\psi\rangle^{A_{1}A_{2}\cdots A_{k}}|\psi\rangle^{A_{k+1}}\cdots|\psi\rangle^{A_{n}})\geqslant E_{(k)}(\rho^{A_{1}A_{2}\cdots A_{k-1}}\otimes|\psi\rangle\langle\psi|^{A_{k+1}}\cdots|\psi\rangle\langle\psi|^{A_{n}})=0$ for any $|\psi\rangle^{A_{1}A_{2}\cdots A_{k}}|\psi\rangle^{A_{k+1}}\cdots|\psi\rangle^{A_{n}}$, $\rho^{A_{1}A_{2}\cdots A_{k-1}}={\rm Tr}_{A_{k}}|\psi\rangle\langle\psi|^{A_{1}A_{2}\cdots A_{k}}$. Therefore item (iv) is straightforward from this point of view. Hereafter, if a $k$-PEM $E_{(k)}$ obeys Eq. (11) and Eq. (12), we call it is $k$-monotonic and coarsening monotonic, respectively.
@@
which is larger than $E_{(3)}(|\psi\rangle)$ whenever $|\psi\rangle^{DE}$, $|\psi\rangle^{GH}$ and $|\psi\rangle^{IJ}$ are entangled states. Namely, Eq. (13) is violated under the general coarsening operation of type (b). We thus adjust the hierarchy condition (i.e., the tight coarsening monotone condition) as Eq. (13) in stead of the “strict” hierarchy condition in Refs. Guo2024pra ; Guo2020pra .
@@
for any genuinely entangled pure state $|\psi\rangle\in\mathcal{H}^{ABCD}$.
@@ 4 Two classes of $k$-PEMos
In this section, we give two classes of $k$-PEMos, where the first class is based on the unified multipartite entanglement measure introduced in Refs. Guo2020pra ; Guo2022jpa ; Guo2024rip and the second class is similar to that of the $k$-entanglement measure defined by the minimal sum of the reduced functions in Ref. Guo2024pra . Hereafter, if a measure $E$ of entanglement for bipartite pure state is defined via some function of the reduced state, i.e., $E(|\psi\rangle^{AB})=h(\rho^{A})$, such a function $h$ is called the reduced function of $E$ Guo2024pra . For example, the concurrence of $|\psi\rangle^{AB}$ is defined by $C(|\psi\rangle^{AB})=\sqrt{2(1-{\rm Tr}\rho_{A}^{2})}$, then $h_{C}(\rho)=\sqrt{2(1-{\rm Tr}\rho^{2})}$ is the reduced function of $C$.
@@ 4.1 $k$-PEMo from unified multipartite entanglement measure
If $|\psi\rangle=|\psi\rangle^{AB}|\psi\rangle^{CDE}|\psi\rangle^{FGH}|\psi\rangle^{I}$ with $|\psi\rangle^{CDE}$ and $|\psi\rangle^{FGH}$ are genuinely entangled, then the 3-partite entanglement is only contained in $|\psi\rangle^{CDE}$ and $|\psi\rangle^{FGH}$. The 3-partite entanglement of $|\psi\rangle$ can be quantified as $E^{(3)}(|\psi\rangle^{CDE})+E^{(3)}(|\psi\rangle^{FGH})$ for some unified multipartite entanglement measure $E^{(k)}$ (the unified multipartite entanglement measure (MEM) was introduced in Ref. Guo2020pra ; Guo2024rip , e.g., $E^{(n)}(|\psi\rangle^{A_{1}A_{2}\cdots A_{n}})=\frac{1}{2}\sum_{i}S(\rho^{A_{i}})$ is a unified MEM, where $S(\rho)=-{\rm Tr}(\rho\log_{2}\rho)$ is the von Neumann entropy of $\rho$). Similarly, the 2-partite entanglement should be $E^{(2)}(\psi\rangle^{AB})+E^{(3)}(|\psi\rangle^{CDE})+E^{(3)}(|\psi\rangle^{FGH})$.
@@
Proof
Items (i) and (ii) are straightforward. For any $|\psi\rangle\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$, we suppose that $E_{(k)}(|\psi\rangle)=\sum\limits_{j=1}^{r}E^{(s(t_{j}))}(|\psi\rangle^{X_{t_{j}}})>0$ for some partition $X_{t_{1}}|X_{t_{2}}|\cdots|X_{t_{r}}|X_{*}$. It turns out that $E_{(k)}(|\psi\rangle^{X_{t_{j}}})=E_{(k-1)}(|\psi\rangle^{X_{t_{j}}})$ for any $j$, $E_{(k)}(|\phi\rangle^{X_{*}})=0$, but it is possible that $E_{(k-1)}(|\phi\rangle^{X_{*}})>0$, which implies (iii) is true. For any partition $X_{1}|X_{2}|\cdots|X_{p}$ of $A_{1}A_{2}\cdots A_{n}$, $p\geqslant k$, we consider $X_{2}|\cdots|X_{p}$ w.n.l.g., namely the partition that by discarding $X_{1}$ from $X_{1}|X_{2}|\cdots|X_{p}$. If $X_{1}$ is $X_{t_{j}}$ or some subsytem(ssome) of $X_{t_{j}}$, (iv) is clear since $E^{(s)}$ is unified Guo2024rip [a unified MEM is decreasing under the coarsening relation of type (a)]. If $X_{1}$ is $X_{*}$ or some subsytem(s) of $X_{*}$, (iv) is clear since $E_{(k)}(|\psi\rangle)=E_{(k)}(\rho^{X_{2}|\cdots|X_{p}})$. If $X_{1}$ cotains some subsytem(s) of some $X_{t_{j}}$, (iv) is also ture since $E^{(s)}$ is unified. The other cases can be argued similarly.
If the reduced function $h$ is subadditive, then $E^{(n)}$ is complete Guo2020pra ; Guo2024rip , which means that $E^{(n)}$ is nonincreasing under the coarsening operation of type (b). This completes the proof.
Another candidate for the unified global MEM is the one defined by the sum of all bipartite entanglement Guo2022jpa , i.e.,
@@
where $h$ is some given reduced function. We denote by $\mathcal{E}_{(k)}$ the quantity that is defined as in Eq. (20) just with $\mathcal{E}^{(s(t_{j}))}$ replacing $E^{(s(t_{j}))}$. Using similar arguments as in the proof of Theorem 4.1, we can conclude the following theorem.
Theorem 4.2
@@
In Eq. (20), $|\psi\rangle^{X_{t_{j}}}$ is genuinely entangled, so $E^{(s(t_{j}))}$ can be any genuine entanglement measure instead. For example, the genuine entanglement measure from the minimal reduced function, which is defined by Guo2024rip
$\varepsilon_{g^{\prime\prime}}^{(n)}(|\psi\rangle^{A_{1}A_{2}\cdots A_{n}})=\min\limits_{X}h(\rho^{X})$
where $h$ is some given reduced function, $X\subsetneq\{A_{1},A_{2},\dots,A_{n}\}$, namely the minimum runs over all possible reduced states. Then the corresponding $k$-PEMo is not unified in general since $\varepsilon_{g^{\prime\prime}}^{(n)}$ may increase under the coarsening relation of type (a) Guo2024rip .